# NoProp: Training Neural Networks without Back-propagation or Forward-propagation

### v1

@@ Abstract
The canonical deep learning approach for learning requires computing a gradient term at each layer by back-propagating the error signal from the output towards each learnable parameter. Given the stacked structure of neural networks, where each layer builds on the representation of the layer below, this approach leads to hierarchical representations. More abstract features live on the top layers of the model, while features on lower layers are expected to be less abstract. In contrast to this, we introduce a new learning method named NoProp, which does not rely on either forward or backwards propagation. Instead, NoProp takes inspiration from diffusion and flow matching methods, where each layer independently learns to denoise a noisy target. We believe this work takes a first step towards introducing a new family of gradient-free learning methods, that does not learn hierarchical representations – at least not in the usual sense. NoProp needs to fix the representation at each layer beforehand to a noised version of the target, learning a local denoising process that can then be exploited at inference. We demonstrate the effectiveness of our method on MNIST, CIFAR-10, and CIFAR-100 image classification benchmarks. Our results show that NoProp is a viable learning algorithm which achieves superior accuracy, is easier to use and computationally more efficient compared to other existing back-propagation-free methods. By departing from the traditional gradient based learning paradigm, NoProp alters how credit assignment is done within the network, enabling more efficient distributed learning as well as potentially impacting other characteristics of the learning process.
@@ 1 Introduction
Back-propagation (Rumelhart et al., 1986) has long been a cornerstone of deep learning, and its application has enabled deep learning technologies to achieve remarkable successes across a wide range of domains from sciences to industries. Briefly, it is an iterative algorithm, which adapts the parameters of a multi-layer neural network at each step such that its output match better a desired target. Each step of back-propagation first performs a forward-propagation of the input signal to generate a prediction, then compares the prediction to a desired target, and finally propagates the error signal back through the network to determine how the weights of each layer should be adjusted to decrease the error. This way, each layer can be thought of as learning to change the representation it receives from the layer below it to one that subsequent layers can use to make better predictions. The error signal propagated backwards is used to do credit assignment, i.e. to decide how much each parameter needs to change in order to minimize the error.
@@
Alternative optimization methods to back-propagation include gradient-free methods (e.g., direct search methods (Fermi, 1952; Torczon, 1991) and model-based methods (Bortz & Kelley, 1997; Conn et al., 2000)), zero-order gradient methods (Flaxman et al., 2004; Duchi et al., 2015; Nesterov & Spokoiny, 2015; Liu et al., 2020; Ren et al., 2022) that use finite difference to approximate gradients, evolution strategies (Wierstra et al., 2014; Salimans et al., 2017; Such et al., 2017; Khadka & Tumer, 2018), methods that rely on local losses such as difference target propagation (Lee et al., 2015) and the forward-forward algorithm (Hinton, 2022). However, these methods have not been widely adopted for learning neural networks due to their limitations in terms of accuracy, computational efficiency, reliability and scalability.
In this paper, we propose a back-propagation-free approach for training. The method is based on the denoising score matching approach that underlies diffusion models (Sohl-Dickstein et al., 2015; Chen et al., 2018; Ho et al., 2020; Song et al., 2021), enabling each layer of the neural network to be trained independently. In brief, at training time each layer is trained to predict the target label given a noisy label and the training input, while at inference time each layer takes the noisy label produced by the previous layer, and denoises it by taking a step towards the label it predicts. Of particular note is the observation that at training time the method does not even require a forward pass, hence we call our method NoProp. We show experimentally on the MNIST, CIFAR-10 and CIFAR-100 benchmarks that NoProp is significantly more performant than previous back-propagation-free methods while at the same time being simpler, more robust, and more computationally efficient.
@@ 2 Methodology
In this section we will describe the NoProp method for back-propagation-free learning. While the technical ideas follow closely those in variational diffusion models (Sohl-Dickstein et al., 2015; Kingma et al., 2021; Gulrajani & Hashimoto, 2024), we apply them in a different context and interpret them differently.
@@ 2.1 NoProp
Figure 1: Architecture of NoProp. $z_{0}$ represents Gaussian noise, while $z_{1},\dots,z_{T}$ are successive transformations of $z_{0}$ through the learned dynamics $u_{1},\dots,u_{T}$, with each layer conditioned on the image $x$, ultimately producing the class prediction $\hat{y}$.
@@
We see that each $\hat{u}_{\theta_{t}}(z_{t-1},x)$ is trained to directly predict $u_{y}$ given $z_{t-1}$ and $x$ with an L2 loss, while $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$ is trained to minimise cross-entropy loss. Each block $\hat{u}_{\theta_{t}}(z_{t-1},x)$ is trained independently, and this is achieved without forward- or back-propagation across the network.
@@ 2.2.2 Continuous-time diffusion models and neural ODEs
The NoProp framework presented above involves learning $\hat{u}_{\theta_{t}}(z_{t-1},x)$ parameterised by $\theta_{t}$ for each layer $t=1,2,\ldots,T$ along with a linear layer for $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$. We can extend this to a continuous version where the number of latent variables $T$ tends to infinity (Kingma et al., 2021), leading to the objective
@@ 2.2.3 Flow matching
where $v_{\theta}(z_{t},x,t)$ is a neural network with parameters $\theta$. When the label embeddings are jointly learned, we introduce an additional anchor loss to prevent different class embeddings from collapsing. Following prior work (Gao et al., 2022; Hu et al., 2024), we incorporate a cross-entropy term $-\log\hat{p}_{\theta_{\mathrm{out}}}(y\mid\tilde{z}_{1}(z_{t},x,t))$, where $\tilde{z}_{1}(z_{t},x,t)$ is an extrapolated linear estimate
@@ 2.3.2 Training procedure
The models used for training are illustrated in Figure 6 in the Appendix. For discrete-time diffusion, we use a neural network $\hat{u}_{\theta_{t}}$ to model the diffusion dynamics $u_{t}$. The model takes an input image $x$ and a latent variable $z_{t-1}$ in the label embedding space, processing them through separate embedding pathways before concatenation. The image $x$ is passed through a convolutional embedding module followed by a fully connected layer. When the embedding dimension matches the image dimension, $z_{t-1}$ is treated as an image and embedded similarly using a convolutional module. Otherwise, $z_{t-1}$ is passed through a fully connected network with skip connections. The fused representation is then processed by additional fully connected layers, producing logits as output. The diffusion dynamics $\hat{u}_{\theta_{t}}$ are obtained by applying a softmax function to the logits, yielding a probability distribution over class embeddings. The final output $\hat{u}_{\theta_{t}}$ is then computed as a weighted sum of the class embeddings in $W_{\mathrm{Embed}}$ using this probability distribution.
@@ 3 Related Works
3.1 Back-propagation-Free Methods
Forward-forward algorithm (Hinton, 2022) replaces the traditional forward and backward passes with two forward passes in image classification: one using images paired with correct labels and the other with images paired with incorrect labels.
Target propagation and its variant (e.g. Lee et al., 2015; Lillicrap et al., 2016) work by learning an inverse mapping from the output layer to the hidden layers, producing local targets for weight updates without back-propagation. However, their performance heavily depends on the quality of the learned inverse mappings, and inaccuracies in the autoencoders can lead to suboptimal updates, limiting their effectiveness in deep networks.
Zero-order methods (Flaxman et al., 2004; Duchi et al., 2015; Ren et al., 2022) is a broad family of algorithms that estimate gradients without explicit differentiation, making them useful for black-box and non-differentiable functions.
Evolution strategies (Wierstra et al., 2014; Salimans et al., 2017; Khadka & Tumer, 2018), inspired by natural evolution, optimize parameters by perturbing them with random noise and updating in the direction of higher rewards using a population-based optimization approach. However, they require a large number of function evaluations, leading to high sample complexity, and effective exploration remains a key challenge for high-dimensional parameter spaces.
@@ 3.2 Diffusion and flow matching
Several works in diffusion and flow matching are closely related to our method. Han et al. (2022) introduced classification and regression diffusion models, Kim et al. (2025) proposed flow matching for paired data, and Hu et al. (2024) applied flow matching to conditional text generation. In contrast, our paper explores the implications of these ideas within the framework of back-propagation-free learning.
@@ 3.3 Representation Learning
Generally speaking, most alternative methods to back-propagation, whether they simply approximate gradients differently, or utilize a different search scheme — as it is the case of evolution strategies — still rely on learning intermediate representations that build on top of each other (Bengio et al., 2013). This allows learning more abstract representations as we look at deeper layers of the model, and is thought to be of fundamental importance for deep learning and for representing cognitive processes (Markman & Dietrich, 2000). Indeed, initial successes of deep learning have been attributed to the ability to train deep architectures to learn hierarchical representations (Hinton et al., 2006; Bengio et al., 2013), while early interpretability work focused on visualising these increasingly more complex features (Zeiler & Fergus, 2014; Lee et al., 2009).
By getting each layer to learn to denoise labels, with the label noise distribution chosen by the user, we can argue that NoProp does not learn representations. Rather, it relies on representations designed by the user (specifically, the representations in intermediate layers are Gaussian noised embeddings of target labels for diffusion, and an interpolation between Gaussian noise and embeddings of target labels for flow matching). This is perhaps unsurprising: forward- and back-propagation can be understood as information being disseminated across the layers of a neural network in order to enable the representation of each layer to “fit in” with those of neighbouring layers and such that the target label can be easily predicted from the representation at the last layer. So for NoProp to work without forward- or back-propagation, these layer representations have to be fixed beforehand, i.e. to be designed by the user.
The fact that NoProp achieves good performance without representation learning leads to the question of whether representation learning is in fact necessary for deep learning, and whether by designing representations we can enable alternative approaches to deep learning with different characteristics. Specifically, note that representations fixed in NoProp are not those one might think of as being more abstract in later layers, which opens the door to revisiting the role of hierarchical representations in modelling complex behaviour (Markman & Dietrich, 2000). These questions can become increasingly important as core assumptions of back-propagation based learning, like the i.i.d. assumption and sequential propagation of information and error signal through the network, are proving to be limiting.
@@ 4 Experiments
Datasets.
We use three benchmark datasets: MNIST, CIFAR-10, and CIFAR-100. The MNIST dataset consists of 70,000 grayscale images of handwritten digits (28x28 pixels), spanning 10 classes (digits 0-9). CIFAR-10 contains 60,000 color images (32x32 pixels) across 10 different object classes. CIFAR-100, similar in structure to CIFAR-10, includes 60,000 color images, but it is divided into 100 fine-grained classes. We use the standard training/test set splits. In our experiments, we do not apply any data augmentation techniques to these datasets.
@@
We refer to this method as NoProp with Discrete-Time Diffusion (NoProp-DT). We fix $T=10$ and, during each epoch, update the parameters for each time step sequentially, as outlined in Algorithm 1. While sequential updates are not strictly necessary for the algorithm to work — as time steps can also be sampled — we choose this approach to align with other back-propagation-free methods, ensuring consistency with existing approaches in the literature. In terms of parameterisation of the class probabilities $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$ in Equation 8, we explore two approaches. The first, as described earlier, involves using softmax on the outputs of a fully connected layer $f$ with parameters $\theta_{\textrm{out}}$, such that
@@
where $\alpha_{1},\dots,\alpha_{T}$ are learnable parameters constrained to the range $(-1,1)$, defined as $\alpha_{t}=\tanh(w_{t})$ with learnable $w_{1},\dots,w_{T}$. This design closely resembles the forward pass of NoProp-DT, but with the network trained by standard back-propagation. The $\hat{u}_{\theta_{t}}(z_{t-1},x)$ and $\hat{p}_{\theta_{\mathrm{out}}}(y|z_{T})$ have identical model structures to those in NoProp-DT.
@@
Previous back-propagation-free methods
@@
Our results, summarised in Table 1, demonstrate that NoProp-DT achieves performance comparable to or better than back-propagation on MNIST, CIFAR-10, and CIFAR-100 in the discrete-time setting. Moreover, NoProp-DT outperforms prior back-propagation-free methods, including the Forward-Forward Algorithm, Difference Target Propagation (Lee et al., 2015), and a forward gradient method referred to as Local Greedy Forward Gradient Activity-Perturbed (Ren et al., 2022). While these methods use different architectures and do not condition layers explicitly on image inputs as NoProp does—making direct comparisons challenging—our method offers the distinct advantage of not requiring a forward pass. Additionally, NoProp demonstrates reduced GPU memory consumption during training, as shown in Table 2. To illustrate the learned class embeddings, Figure 2 visualises both the initializations and the final class embeddings for CIFAR-10 learned by NoProp-DT, where the embedding dimension matches the image dimension.
@@
We also conducted ablation studies on the parameterisations of class probabilities, $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$, and the initializations of the class embedding matrix, $W_{\text{Embed}}$, with results shown in Figure 4 and Figure 5, respectively. The ablation results reveal no consistent advantage between the class probability parameterisations, with performance varying across datasets. For class embedding initializations, both orthogonal and prototype initializations are generally comparable to or outperform random initializations.

### v2

# NoProp: Training Neural Networks without Full Back-propagation or Full Forward-propagation

@@ Abstract
The canonical deep learning approach for learning requires computing a gradient term at each block by back-propagating the error signal from the output towards each learnable parameter. Given the stacked structure of neural networks, where each block builds on the representation of the block below, this approach leads to hierarchical representations. More abstract features live on the top blocks of the model, while features on lower blocks are expected to be less abstract. In contrast to this, we introduce a new learning method named NoProp, which does not rely on either forward or backwards propagation across the entire network. Instead, NoProp takes inspiration from diffusion and flow matching methods, where each block independently learns to denoise a noisy target using only local targets and back-propagation within the block. We believe this work takes a first step towards introducing a new family of learning methods, that does not learn hierarchical representations – at least not in the usual sense. NoProp needs to fix the representation at each block beforehand to a noised version of the target, learning a local denoising process that can then be exploited at inference. We demonstrate the effectiveness of our method on MNIST, CIFAR-10, and CIFAR-100 image classification benchmarks. Our results show that NoProp is a viable learning algorithm, is easy to use and computationally efficient. By departing from the traditional learning paradigm which requires back-propagating a global error signal, NoProp alters how credit assignment is done within the network, enabling more efficient distributed learning as well as potentially impacting other characteristics of the learning process.
@@ 1 Introduction
Back-propagation (Rumelhart et al., 1986) has long been a cornerstone of deep learning, and its application has enabled deep learning technologies to achieve remarkable successes across a wide range of domains from sciences to industries. Briefly, it is an iterative algorithm, which adapts the parameters of a multi-block neural network at each step such that its output match better a desired target. In this paper, we define a block as either a single layer or a group of consecutive layers within a neural network. Each step of back-propagation first performs a forward-propagation of the input signal to generate a prediction, then compares the prediction to a desired target, and finally propagates the error signal back through the network to determine how the weights of each block should be adjusted to decrease the error. This way, each block can be thought of as learning to change the representation it receives from the block below it to one that subsequent blocks can use to make better predictions. The error signal propagated backwards is used to do credit assignment, i.e. to decide how much each parameter needs to change in order to minimize the error.
@@
Alternative optimization methods to back-propagation include gradient-free methods (e.g., direct search methods (Fermi, 1952; Torczon, 1991) and model-based methods (Bortz & Kelley, 1997; Conn et al., 2000)), zero-order gradient methods (Flaxman et al., 2004; Duchi et al., 2015; Nesterov & Spokoiny, 2015; Liu et al., 2020; Ren et al., 2022), evolution strategies (Wierstra et al., 2014; Salimans et al., 2017; Such et al., 2017; Khadka & Tumer, 2018), methods that rely on local losses such as difference target propagation (Lee et al., 2015) and the forward-forward algorithm (Hinton, 2022). However, such approaches often retain standard neural network architectures and seek only to modify the optimization algorithm, and frequently fall short of back-propagation in terms of accuracy, computational efficiency and scalability.
In this paper, we explore an alternative path, rethinking both the architecture and the training process. We propose a novel training framework that avoids global back-propagation across network blocks, based on the denoising score matching approach that underlies diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021). Unlike prior methods that approximate or replace back-propagation within standard architectures, NoProp employs a block-wise design where inputs are broadcast to each block, allowing each block to be trained independently using only local back-propagation. In brief, at training time each block is trained to predict the target label given a noisy label and the training input using back-propagation within the block, while at inference time each block takes the noisy label produced by the previous block, and denoises it by taking a step towards the label it predicts. This design avoids full forward and backward passes across the entire network, enabling block-local updates with minimal coordination, such as shared embeddings. While NoProp still uses back-propagation within blocks, its local training structure makes it more biologically plausible than full back-propagation and naturally suited for parallel and distributed training. We evaluate NoProp on MNIST, CIFAR-10, and CIFAR-100, demonstrating performance comparable to full back-propagation within the architecture we study. We also show that NoProp achieves better performance and efficiency than the adjoint sensitivity method used for training neural ordinary differential equations (Neural ODEs) (Chen et al., 2018). While direct comparisons to other methods are limited by architectural differences, our aim is to show that local learning strategies inspired by generative models can be effectively applied to supervised learning. Although we have not fully explored scaling NoProp to larger datasets or more complex domains, we view NoProp as an early step toward more efficient, biologically-inspired, modular, and parallelizable alternatives to end-to-end back-propagation.
@@ 2 Methodology
In this section we will describe the NoProp method for learning without full forward- or back-propagation. While the technical ideas are an application of variational diffusion models (Sohl-Dickstein et al., 2015; Kingma et al., 2021; Gulrajani & Hashimoto, 2024), we apply them in the context of alternative methods to end-to-end back-propagation, and interpret them differently as enabling learning without full forward or backward passes.
@@ 2.1 NoProp
Figure 1: Architecture of NoProp. $z_{0}$ represents Gaussian noise, while $z_{1},\dots,z_{T}$ are successive transformations of $z_{0}$ through the learned dynamics $u_{1},\dots,u_{T}$, with each block conditioned on the image $x$, ultimately producing the class prediction $\hat{y}$.
@@
We see that each $\hat{u}_{\theta_{t}}(z_{t-1},x)$ is trained to directly predict $u_{y}$ given $z_{t-1}$ and $x$ with an L2 loss, while $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$ is trained to minimise cross-entropy loss. Each block $\hat{u}_{\theta_{t}}(z_{t-1},x)$ is trained independently, and this is achieved without full forward- or back-propagation across the entire network.
@@ 2.2.2 Continuous-time diffusion models and neural ODEs
The NoProp framework presented above involves learning $\hat{u}_{\theta_{t}}(z_{t-1},x)$ parameterised by $\theta_{t}$ for each time step $t=1,2,\ldots,T$ along with a linear layer for $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$. We can extend this to a continuous version where the number of latent variables $T$ tends to infinity (Kingma et al., 2021), leading to the objective
@@ 2.2.3 Flow matching
where $v_{\theta}(z_{t},x,t)$ is a neural network with parameters $\theta$. When the label embeddings are jointly learned, we introduce an additional anchor loss to prevent different class embeddings from collapsing. Following prior work (Gao et al., 2022; Hu et al., 2024), we incorporate a cross-entropy term $\log\hat{p}_{\theta_{\mathrm{out}}}(y\mid\tilde{z}_{1}(z_{t},x,t))$, where $\tilde{z}_{1}(z_{t},x,t)$ is an extrapolated linear estimate
@@ 2.3.2 Training procedure
The models used for training are illustrated in Figure 5 in the Appendix. For discrete-time diffusion, we use a neural network $\hat{u}_{\theta_{t}}$ to model the diffusion dynamics $u_{t}$. The model takes an input image $x$ and a latent variable $z_{t-1}$ in the label embedding space, processing them through separate embedding pathways before concatenation. The image $x$ is passed through a convolutional embedding module followed by a fully connected layer. When the embedding dimension matches the image dimension, $z_{t-1}$ is treated as an image and embedded similarly using a convolutional module. Otherwise, $z_{t-1}$ is passed through a fully connected network with skip connections. The fused representation is then processed by additional fully connected layers, producing logits. We then apply a softmax function to the logits, yielding a probability distribution over class embeddings. The final output of $\hat{u}_{\theta_{t}}$ is then computed as a weighted sum of the class embeddings in $W_{\mathrm{Embed}}$ using this probability distribution.
@@ 3 Related Works
3.1 Alternative Methods to Back-propagation
Figure 2: Comparison of weight update strategies in end-to-end back-propagation and its alternatives. a Feedback Alignment closely resembles standard end-to-end back-propagation. b A variety of optimization methods that eliminate the need for backward passes entirely. c After a forward pass, each block updates its weights using local unsupervised or self-supervised objectives, while previous blocks are kept frozen. d Similar to c, but the target labels are broadcast to each block, enabling local supervised updates. e NoProp broadcasts both the input and the labels to every block, allowing each to update independently using only local back-propagation.
A wide range of alternatives to back-propagation have been proposed, aiming either to improve biological plausibility, increase computational efficiency, or both. Below, we highlight key approahces and illustrate their weight updates in Figure 2. One well-known limitation of backpropagation is the “weight transport problem” (Grossberg, 1987), which refers to the biologically implausible requirement that the backward pass must have access to the exact weights used in the forward pass. Feedback Alignment (Lillicrap et al., 2016) addresses this by replacing the backward weights with fixed, random matrices. Although the network still relies on a top-down propagation of errors, this method avoids symmetric weight transport, making it more biologically plausible. However, it does not reduce computational cost since it still requires sequential backward passes.
Several optimization approaches avoid backward passes entirely by using only forward computations to update parameters. These include zeroth-order gradient estimators (Flaxman et al., 2004; Duchi et al., 2015; Nesterov & Spokoiny, 2015; Liu et al., 2020) and evolutionary strategies (Wierstra et al., 2014; Salimans et al., 2017; Khadka & Tumer, 2018). Evolution strategies are inspired by natural selection and operate without gradient information, making them well-suited for black-box optimization. However, they are typically sample-inefficient, especially in high-dimensional spaces, and require large numbers of function evaluations. A related but more efficient alternative is forward gradient descent (Baydin et al., 2022), which uses only forward passes to compute an unbiased estimate of the gradient and reduces computational cost compared to back-propagation.
Another prominent class of methods focuses on block-wise local learning, where the network is divided into modules or blocks that are trained independently using local objectives. These methods eliminate the need for error signals to propagate through the entire network and are naturally amenable to parallel training. Early examples of this idea appear in deep belief networks (Hinton et al., 2006). In more recent work, Direct Feedback Alignment (Nøkland, 2016) extends the original Feedback Alignment idea by applying independent random feedback paths to each block. Greedy layerwise training with auxiliary networks (Belilovsky et al., 2019) appends auxiliary networks to each block to optimize a supervised learning objective for that block. Clark et al. (2021) propose broadcasting a global error signal to all hidden units, eliminating the need for unit-specific feedback. The Forward-Forward algorithm (Hinton, 2022) trains each block by contrasting its response to positive and negative examples.
An alternative branch of block-wise methods relies on unsupervised or self-supervised local objectives. These approaches update each block independently to learn useful representations, which are later used for downstream supervised learning. Examples include Hebbian learning and local learning rules based on reconstruction, contrastive objectives, or predictive coding (Löwe et al., 2019; Illing et al., 2021; Tang et al., 2022; Journé et al., 2022; Halvagal & Zenke, 2023; Siddiqui et al., 2023). These methods maintain biological plausibility through local updates, and they are computationally efficient due to their potential for parallelism.
Several other approaches also decouple updates between layers or blocks. Ren et al. (2022) extend forward gradient descent to a block-wise local training setting. Synthetic gradients (Jaderberg et al., 2017) train small networks to predict the gradient at each layer, allowing for parallel updates. Kickback (Balduzzi et al., 2014) simplifies the gradient computation by retaining only a global error signal. Target propagation (Lee et al., 2015)replaces error back-propagation with a learned inverse function that propagates target activations back through the network. While these methods can increase biological plausibility, their effectiveness often depends on the quality of the learned components, such as the synthetic gradient models or inverse mappings.
Our method, NoProp, builds on the idea of block-wise local learning but differs in key ways. Both the input (e.g., an image in classification tasks) and the label are broadcast to each block of the network. The forward pass is predetermined by a noise schedule, so a full forward pass through the entire network is not required. Each block is trained with a local objective using back-propagation internally, without relying on end-to-end gradient flow. When label embeddings are fixed (e.g., using one-hot vectors), blocks can be updated independently, enabling parallel training and improving computational efficiency relative to standard back-propagation. If the label embeddings or noise schedule are jointly learned with the model, some limited coupling between blocks is introduced, but the method still avoids global back-propagation. In continuous-time settings, NoProp is also more efficient than approaches requiring ODE simulations, such as the adjoint sensitivity method (Chen et al., 2018). Finally, because NoProp does not rely on propagating a global error signal across the network, it offers greater biological plausibility than end-to-end back-propagation.
@@ 3.2 Diffusion and flow matching
Several works in diffusion and flow matching are closely related to our method. Han et al. (2022) introduced classification and regression diffusion models, Kim et al. (2025) proposed flow matching for paired data, and Hu et al. (2024) applied flow matching to conditional text generation. In contrast, our paper explores the implications of these ideas within the framework of alternative learning paradigms, building on the observation that diffusion and flow matching methods do not require full forward and backward passes.
@@ 3.3 Representation Learning
Generally speaking, most alternative methods to back-propagation, whether they simply approximate gradients differently, or utilize a different search scheme — as it is the case of evolution strategies — still rely on learning intermediate representations that build on top of each other (Bengio et al., 2013). This allows learning more abstract representations as we look at deeper blocks of the model, and is thought to be of fundamental importance for deep learning and for representing cognitive processes (Markman & Dietrich, 2000). Indeed, initial successes of deep learning have been attributed to the ability to train deep architectures to learn hierarchical representations (Hinton et al., 2006; Bengio et al., 2013), while early interpretability work focused on visualising these increasingly more complex features (Zeiler & Fergus, 2014; Lee et al., 2009).
By getting each block to learn to denoise labels, with the label noise distribution chosen by the user, we can argue that NoProp does not learn global representations across blocks. Rather, it relies on representations designed by the user (specifically, the representations in intermediate blocks are Gaussian noised embeddings of target labels for diffusion, and an interpolation between Gaussian noise and embeddings of target labels for flow matching). This is perhaps unsurprising: forward- and back-propagation can be understood as information being disseminated across the blocks of a neural network in order to enable the representation of each block to “fit in” with those of neighbouring blocks and such that the target label can be easily predicted from the representation at the last block. So for NoProp to work without forward- or back-propagation, these intermediate representations have to be fixed beforehand, i.e. to be designed by the user.
The fact that NoProp achieves good performance without global representation learning leads to the question of whether such learning is in fact necessary for deep learning, and whether by designing representations we can enable alternative approaches to deep learning with different characteristics. Specifically, note that representations fixed in NoProp are not those one might think of as being more abstract in later blocks, which opens the door to revisiting the role of hierarchical representations in modelling complex behaviour (Markman & Dietrich, 2000). These questions can become increasingly important as core assumptions of back-propagation based learning, like the i.i.d. assumption and sequential propagation of information and error signal through the network, are proving to be limiting.
@@ 4 Experiments
We refer to this method as NoProp with Discrete-Time Diffusion (NoProp-DT). We fix $T=10$ and, during each epoch, update the parameters for each time step sequentially, as outlined in Algorithm 1. While sequential updates are not strictly necessary for the algorithm to work — as time steps can also be sampled — we choose this approach to align with prior alternative methods to full back-propagation, ensuring consistency with existing approaches in the literature. In terms of parameterisation of the class probabilities $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$ in Equation 8, we explore two approaches. The first, as described earlier, involves using softmax on the outputs of a fully connected layer $f$ with parameters $\theta_{\textrm{out}}$, such that
@@
where $\alpha_{1},\dots,\alpha_{T}$ are learnable parameters constrained to the range (1,1)(-1,1), defined as $\alpha_{t}=\tanh(w_{t})$ with learnable $w_{1},\dots,w_{T}$. This design closely resembles the forward pass of NoProp-DT, but with the network trained by standard back-propagation. The $\hat{u}_{\theta_{t}}(z_{t-1},x)$ and $\hat{p}_{\theta_{\mathrm{out}}}(y|z_{T})$ have identical model structures to those in NoProp-DT.
@@
Previous methods without full back-propagation
@@
Method Tiny ImageNet
Train Test
Backprop (one-hot) 75.66±33.67 23.83±6.7
Backprop (prototype) 99.82±0.02 31.26±0.1
NoProp-DT (one-hot) 69.78±1.71 26.55±0.4
NoProp-DT (prototype) 78.45±0.97 25.65±0.19
Table 2: Classification accuracies (%) on Tiny ImageNet in the discrete-time setting. Means and standard errors are computed from 15 values per method, obtained from 3 seeds and 5 inference runs per seed.
@@
Our main results, summarised in Table 1, demonstrate that NoProp-DT achieves performance comparable to or better than back-propagation on MNIST, CIFAR-10, and CIFAR-100 in the discrete-time setting. We also report results from prior alternative methods to full back-propagation, including the Forward-Forward Algorithm, Difference Target Propagation (Lee et al., 2015), and the Local Greedy Forward Gradient Activity-Perturbed method (Ren et al., 2022). However, direct comparisons are challenging because NoProp employs a fundamentally different framework based on conditional diffusion or flow matching, where the input is broadcast to each block (see Figure 2). While our results surpass previously reported performances of these methods, differences in architectures and model sizes make direct comparisons difficult, so improvements cannot be solely credited to our approach. Nevertheless, our results demonstrate that NoProp can achieve non-trivial performance without requiring full forward and backward passes through the entire network. We also conducted a limited evaluation on Tiny ImageNet, comparing NoProp-DT with back-propagation, with results shown in Table 2. While the method has not been fully optimized in this paper and should be viewed as a proof of concept, it provides preliminary evidence that NoProp has the potential to scale beyond small datasets. Additionally, NoProp demonstrates reduced GPU memory consumption during training, as shown in Table 3. To illustrate the learned class embeddings, Figure 3 visualises both the initializations and the final class embeddings for CIFAR-10 learned by NoProp-DT, where the embedding dimension matches the image dimension.
@@
We also conducted ablation studies on the parameterisations of class probabilities, $\hat{p}_{\theta_{\textrm{out}}}(y|z_{T})$, and the initializations of the class embedding matrix, $W_{\text{Embed}}$, with results shown in Figure 6 and Figure 7 in the Appendix. The ablation results reveal no consistent advantage between the class probability parameterisations, with performance varying across datasets. For class embedding initializations, random initializations are comparable to orthogonal and prototype initializations.